OpenStudy (anonymous):

find all numbers for which the rational expression is undefined t^3-4t/t^2-16

OpenStudy (turingtest):

If you're not missing any parentheses then\[t^3-4t/t^2+16=t^3-4/t-16\] is undefined when t=0 because\[0-4/0-16\] is undefined.